3x-5y=18
6x-10y=36

Answer by varralovesjbieber(2)

6x-10y=36
[3x-5y=18]-2
-6x+10y=-36
6x-10y=36
(Cancel out -6x+10y and 6x-10y)
ANSWER: 0=0